Find the area under the normal curve between z = -0.46 and z = 1.55


1
Expert's answer
2021-06-07T13:37:24-0400

"P(-0.46<Z<1.55) = P(Z<1.55) -P(Z< -0.46) \\
= 0.9394 -0.3227 \\
=0.6167"